lynbtsfan03:

write as two terms (divide) 9x-10 / 5

jhonyy9:

do you think it in this way \[\left( \frac{ 9x - 10 }{ 5 } \right) = \frac{ 9x }{ 5 } - \frac{ 10 }{ 5 }\] ?